OpenStudy (anonymous):

If you have a geometric sequence and the "r" value is negative what is the effect? Please help!

OpenStudy (anonymous):

So, I'm not aware of anything "wrong" with a negative r. If you just want to know what it does to the series, we can illustrate that pretty easily. A geometric series is one where we start with a number, a, and then multiply it by a factor, r, for each subsequent term. So the series would be: a, ar, ar^2, ar^3, etc. If r is negative, something will happen with the signs of each term. Assume a is positive. What will the signs of each of the terms be?

OpenStudy (anonymous):

its multiple choice:) a. the # get bigger b. the #s get smaller c. the #s alternate from positive to negative d. the #s don't change. will that help?

OpenStudy (anonymous):

What I'm leading to is one of those choices, so that's good :P

OpenStudy (anonymous):

kay so Im thinkin its a. but im not sure why im thinkin that!?

OpenStudy (anonymous):

Let's say a is positive, and r is negative. Is a*r positive or negative?

OpenStudy (anonymous):

positve

OpenStudy (anonymous):

What's 2* (-3)?

OpenStudy (anonymous):

um.. well that would be -6 right?

OpenStudy (anonymous):

Right. That was a positive number times a negative number. The result is a negative number. So if a is positive, and r is negative, a*r is negative.

OpenStudy (anonymous):

So if a*r is negative, what is a*r^2?

OpenStudy (anonymous):

hold on while I solve it

OpenStudy (anonymous):

-36?

OpenStudy (anonymous):

No. So, when we multiply a positive number with a positive number, we get a positive result. When we multiply a negative number with a negative number, we get a positive result. When we multiply a positive number and a negative number, we get a negative result. So, when we look at the series: a, ar, ar^2, ar^3, etc we see that the signs will alternate if r is negative. r is negative. r^2 = r*r is positive. r^3=r^2 * r is negative. Etc.

OpenStudy (anonymous):

wow. thanks a lot. I never would have figured it out:)